I have a problem here that I have solved, but now I am having a problem graphing it???
I know that the intersection of the lines is (16,63), but I an not understanding how I get TWO lines??? And where to start with each line and how &/or what direction they are going, etc.?? I am lost??
x = person’s forearm
set up:
y = 4x-1 (first piece of rope)
z = 3x +15 (second piece of rope)

4x-1=3x+15
Solve for x
4x=3x+16
1x=16
x=16
Substitute
Y=4x-1
y=4(16)-1
y=64-1
y=63
z=3x+15
z=3(16)+15
z=48+15
z=63

I tried to show my graph, but it is not pasting to this section??

















































Answer by stanbon(75887)

y = 4x-1 (first piece of rope)
z = 3x +15 (second piece of rope)
4x-1=3x+15
-----------
Comment: How can you set 4x-1 equal to 3x+15 when one is "y"
and the other is "z"?
If your variables are really x,y and z you cannot solve the
system of equations with only two equations.
=============
Cheers,
Stan H.




Solve for x
4x=3x+16
1x=16
x=16
Substitute
Y=4x-1
y=4(16)-1
y=64-1
y=63
z=3x+15
z=3(16)+15
z=48+15
z=63